# SpatialLM: Training Large Language Models for Structured Indoor Modeling

### v1

@@ 2.3 Training Schedules
As most recent MLLMs start from pre-trained unimodal backbones, how to align the multimodal content with the language model plays a critical role in the performance of the resulting MLLMs. For example, a common practice adopted by many MLLMs [30, 49] is the inclusion of a two-stage training schedule: (i) an alignment stage which trains the randomly initialized MLP projector only, followed by (ii) a fine-tuning stage that trains both the projector and language model while keeping the vision encoder frozen. Other studies (e.g.[23]) advocate a single-stage training protocol, suggesting that including the projector training stage is unnecessary. Furthermore, when adapting the LLMs to rich visual content such as long videos, some research [68] has shown that a three-stage pipeline, which progressively unfreezes different components in different stages, improves the performance.
In this experiment, we test six possible training configurations, ranging from one-stage to three-stage, to evaluate the influence of different training strategies. As shown in Table 4, we found that a single-stage fine-tuning yields the best results. Adding more stages generally degrades the model performance. We also found that making all parameters trainable in a single stage is important, especially for 3D object detection. This may indicate that, compared to modern image encoders such as SigLIP [61] and DINOv2 [39], current pre-trained point cloud encoders are not as versatile in supporting downstream tasks yet.
3 SpatialLM
In this section, we compare SpatialLM to the current state-of-the-art on scene layout estimation (Section 3.1) and 3D object detection (Section 3.2), respectively. Following our empirical study, we use Qwen2.5-0.5B [42] as the base model and Sonata [56] as the point cloud encoder. We set of resolution at the finest level at $2.5\mathrm{cm}$ and employ a single-stage training on our dataset.
Besides testing on established RGBD benchmarks, we also investigate whether our model can handle inputs from other sources, such as point clouds reconstructed from RGB videos (Section 3.3). Other extensions and more technical details can be found in the appendix.
@@ 3.1 Layout Estimation
Layout estimation focuses on predicting architectural elements, i.e., walls, doors, and windows, within an indoor scene. In this experiment, we include results from two representative baselines, namely RoomFormer [60] and SceneScript [2]. As the state-of-the-art specialist model for layout estimation, RoomFormer employs a highly specialized network architecture, using two-level queries to predict room polygons and their corners, respectively. SceneScript, on the other hand, casts 3D structure estimation as an auto-regressive “next-token-prediction” problem. However, unlike our method, it uses specialized structured language commands to describe the scene and trains a Transformer decoder from scratch.
Following the setting of RoomFormer, we perform evaluation on the Structured3D benchmark [64], which contains 3,500 residential houses with diverse floorplans. We use the original data split of 3000/250/250 for training/validation/testing, respectively. For RoomFormer, we directly download the model checkpoint from its website [^1] https://github.com/ywyue/RoomFormer for inference. Note that RoomFormer takes in 2D density maps projected from point clouds and outputs 2D layouts, which are converted to 3D by extruding based on the lowest/highest points in the point cloud. For SceneScript, we download the pre-trained model and code [^2] https://github.com/facebookresearch/scenescript, and fine-tune it on Structured3D with the default hyperparameter settings.
@@
RoomFormer [60] 70.4 67.2
SceneScript [2] 83.1 80.8
SpatialLM (ft. Structured3D) 32.8 17.9
SpatialLM (ft. Ours) 51.2 38.3
SpatialLM (ft. Ours $\rightarrow$ Structured3D) 86.5 84.6
For our model, we include three variants: (i) trained on Structured3D only, (ii) trained on our dataset only, and (iii) first trained on our dataset, then fine-tuned on Structured3D. As can be seen in Table 5, directly training LLMs on the smaller Structured3D dataset leads to poor performance. In contrast, by first training on our new dataset and then on Structured3D, our model outperforms both baselines for this task. This demonstrates the benefits of our new dataset.
@@
Ground truth RoomFormer [60] SceneScript [2] SpatialLM
@@
Figure 4 provides some qualitative comparisons. Since RoomFormer detects room elements independently, we can see that some doors and windows are not attached to any wall. Auto-regressive methods like SceneScript and our method can preserve the relationships between structures by design. But we observe that SceneScript tends to miss some elements in its output.
@@
Next, we compare SpatialLM to existing methods on 3D object detection. We include results from two representative baselines, namely V-DETR [47] and SceneScript. As the state-of-the-art specialist model for the task, V-DETR builds upon the DETR framework [5] and makes several task-specific improvements, including a 3D vertex relative position encoding (3DV-RPE), object-based normalization, among others. Following prior practice, we perform evaluation on ScanNet [13], which contains 1,513 3D indoor scans with annotations of 18 object categories. The training set is composed of 1,201 scenes, while 312 scenes are used for validation/testing. Since auto-regressive models (i.e., SpatialLM and SceneScript) do not produce confidence scores, we report F1 scores instead of mean Average Precision (mAP), as suggested in [2].
@@
V-DETR [47] 65.1 56.8
SceneScript [2] 49.1 36.8
SpatialLM (ft. ScanNet) 2.9 0.7
SpatialLM (ft. Ours) 33.8 22.6
SpatialLM (ft. Ours $\rightarrow$ ScanNet) 65.6 52.6
For V-DETR, we directly use the model checkpoint downloaded from its website [^3] https://github.com/V-DETR/V-DETR. A confidence threshold of 0.5 is applied for both objectness and semantic predictions, followed by non-maximum suppression (NMS). For SceneScript, we fine-tune the model on ScanNet with the default hyperparameter settings. During fine-tuning, we apply the same data augmentations as used in PTv3 [55] and V-DETR [47].
For our model, we again include three variants: (i) trained on ScanNet only, (ii) trained on our dataset only, and (iii) first trained on our dataset, then fine-tuned on ScanNet. As can be seen in Table 6, the first variant has very low F1 scores, suggesting that ScanNet itself is too small to train LLMs on. By first training on our new dataset and then on ScanNet, our model outperforms SceneScript and achieves competitive results against the state-of-the-art specialist V-DETR.
@@
Ground truth V-DETR [47] SceneScript [2] SpatialLM
@@
Figure 5 shows some qualitative results. Compared to SceneScript, SpatialLM demonstrates improved detection accuracies across categories. In comparison to V-DETR, we observe that our method has some difficulty in locating objects in “picture”, “sink” and “showercurtain” classes. Notably, “picture” and “sink” are the smallest objects in ScanNet by volume ($<0.1\mathrm{m^{3}}$), whereas “showercurtain” category is not included in our dataset and has the fewest occurrences in the ScanNet.
@@ 3.3 Zero-shot Detection on Videos
Although significant progress has been made in SFM and SLAM techniques over the past decades, reliably inferring 3D structure from videos remains a highly challenging problem. Recently, DUSt3R [52] and its extensions [27, 15, 36, 31] provide a novel solution which employs deep networks for simultaneous point matching and camera motion estimation. These techniques are shown to be remarkably robust in dealing with real-world videos or unconstrained image collections.
In this experiment, we collect a set of 107 videos of virtual indoor scene tours and use MASt3R-SLAM [36] to reconstruct a point cloud from each video. Then, we test SpatialLM in a zero-shot setting on the reconstructed point clouds. Unlike RGBD scans, these reconstructions have significant noises, occlusions, and geometric artifacts, posing a greater challenge for structured 3D understanding.
Figure 6 shows some qualitative results. Without fine-tuning, SpatialLM demonstrates strong resilience to imperfect data and preserves consistent layout and object-level predictions. Trained on synthetic datasets with accurate bounding boxes, the model infers full object sizes and orientations from sparse or occluded inputs. In the first example, it predicts beds and nightstands extending to the floor, even when those regions are not visible. Leveraging the auto-regressive nature of LLMs, SpatialLM is also capable of generating plausible room layouts from partial observations. In the second example, the model reasonably reconstructs areas such as the balcony and dining space, filling in missing elements based on contextual understanding.
4 Related Work
4.1 Structured Indoor Modeling
In the literature, some studies focus on room layout estimation from point clouds. Early methods [37, 38] rely on traditional geometric analysis algorithms (i.e., RANSAC) to detect wall planes. SceneCAD [1] jointly optimizes the layout estimation and CAD model alignment of the scanned scene. RoomFormer [60] proposes to train a two-stage Transformer-based network to predict corners and rooms. Meanwhile, a number of recent methods [35, 46, 51, 47, 25] have advanced the state-of-the-art on 3D object detection from point clouds.
@@ 4.1 Structured Indoor Modeling
The closest work to ours is SceneScript [2], which also casts structured indoor reconstruction as a sequence modeling problem. It demonstrates great flexibility as the language can be tailored to different scenarios without changing the method. But unlike our method, SceneScript uses domain-specific tokens for language commands thus requires a specialized decoder.
4.2 Large Language Models for 3D Scene Understanding
Recently, there has been a growing interest in empowering LLMs to understand and reason about 3D spaces, as defined in public benchmarks such as ScanRefer [9] and ScanQA [3]. One line of work [53, 20, 19, 29, 22] first uses 3D detectors to locate objects in the point cloud, then extracts features of individual objects using a pre-trained object-level encoder. In this paper, we study how such 3D detectors can be learned with modern LLMs.
Other methods directly extract scene-level features from the point cloud and align the features with LLMs [18, 66, 10, 16]. 3D-LLM [18] uses pre-trained image encoder to obtain 2D features from multi-view images and back-projects onto the point cloud [21]. It then leverages a Q-former [28] to map them to a fixed number of tokens as input to the LLM. LL3DA [10] adopts a 3D visual encoder pre-trained on ScanNet detection as the scene encoder. ScanReason [66] further employs a 3D point cloud encoder to obtain fine-grained geometric information for accurate object grounding. Scene-LLM [16] tokenizes the 3D point cloud features by simply dividing the space into a fixed-resolution voxel grid and aggregating features in the same voxel.
Besides point cloud-based representations, recent studies have explored other ways to integrate 3D information in LLMs. SpatialVLM [8] primarily focuses on synthesizing large-scale 3D VQA data to enhance the spatial reasoning capabilities of the VLM. LLaVA-3D [67] injects 3D position embeddings into 2D CLIP features to establish a unified architecture for 2D image understanding and 3D scene understanding. SpatialRGPT [11] extracts a 3D scene graph from a single image to enhance grounded spatial reasoning in LLMs. Video-3D LLM [63] learns representations from video frames and 3D coordinates to leverage the strong priors of pre-trained 2D Video LLMs. LSceneLLM [65] uses an adaptive self-attention module to capture fine-grained details in local areas.
@@ 5 Conclusion
In this study, we demonstrate the feasibility of training LLMs for structured indoor modeling tasks. We regard it as a meaningful step towards building future foundation models which can not only understand, but also reason about, interact with, and even create structured 3D scenes.
Limitations. SpatialLM has several limitations. First, it falls short of delivering a universal, state-of-the-art model for extracting 3D structure from arbitrary point clouds. There are significant differences between point clouds from different sources, such as monocular videos, RGBD scans, and LiDAR sensors. As shown in Section 3, although our model exhibits reasonable generalizability across datasets, fine-tuning on a specific dataset is still needed to achieve the best performance. Further scaling up the data and model size could be a promising future direction. Second, in this paper we focus on training LLM for structured indoor modeling only, without considering how it affects the model’s natural skills in natural language processing and reasoning. To address this, a more thorough benchmarking may be performed. Third, our current approach models indoor layouts using a set of predefined object categories, limiting the LLMs in leveraging their open-ended language capabilities. Future directions include extending our work to support open-vocabulary object detection and 3D visual question answering (VQA), enabling more flexible and generalizable scene understanding.
@@ Appendix A SpatialLM Dataset
Object categories. We have curated a set of 59 commonly occurring object categories, organized by functional and semantic similarity, as shown in Table 7. Statistics of these categories can be found in Figure 7.
@@
Figure 8 plots the object-to-room correlation heatmap, which reveals strong correlations between object distribution and specific room types. Additionally, each room type contains a variety of object categories, highlighting the diversity of our dataset.
@@
Dataset visualizations. We provide visualizations of the ground-truth point cloud and layout annotations in Figure 9 to showcase the quality of the photo-realistic scans and the realism of the layout. Additionally, we show rendered images of objects in Figure 10 to illustrate the diversity and quality of the objects in our dataset.
@@ Appendix B Implementation Details of SpatialLM
In Figure 11, we show the full prompt and an example response of SpatialLM. Note that we shift the point cloud and the corresponding layout and object box coordinates to ensure that they are all non-negative. Next, we quantize the coordinates into 1,280 bins, each with a resolution of $2.5\mathrm{cm}$. LLM predicts integer values, which are mapped back to continuous values using the inverse transformation and normalization, restoring them to the original coordinate system.
Data augmentations. We apply a number of augmentations to the point cloud, which are summarized in Table 8 and explained below.
We begin with a cuboid crop following V-DETR [47]. This process involves randomly selecting a point as the center of the cuboid and then determining random dimensions for the cuboid relative to the input point cloud size. We maintain a minimum aspect ratio of 0.8 and ensure that at least 50000 points remain after the cropping process.
@@
cuboid crop min_points=50,000, aspect=0.8, min_crop=0.75, max_crop=1.0, $p$=0.1
random jitter $\sigma$=0.025, clip=0.05, ratio=0.8, $p$=0.9
random jitter $\sigma$=0.2, clip=0.2, ratio=0.05, $p$=0.8
random jitter $\sigma$=0.4, clip=1.0, ratio=0.001, $p$=0.75
random jitter $\sigma$=0.5, clip=4.0, ratio=0.0005, $p$=0.65
elastic distort params=[[0.2, 0.4], [0.8, 1.6]], $p$=[0.8, 0.5]
@@
auto contrast $p$=0.2
chromatic translation $p$=0.75, ratio=0.1
chromatic jitter std=0.05; $p$=0.8
color drop $p$=0.1
@@
For experiments on Structured3D [64] and ScanNet [13], we use a batch size of 8. Fine-tuning on Structured3D involves 50 epochs, while keeping the other settings the same. For fine-tuning on ScanNet, we introduce an additional pre-training stage to mitigate the difference in label spaces across datasets. Specifically, we map the classes in our dataset to those of ScanNet, and further train the model on our dataset for 3 epochs, followed by fine-tuning on ScanNet for 30 epochs.

### v2

@@ 3.3 Training Schedules
$\circ\bullet\bullet$ - - 73.8 68.4 40.2 22.7
$\bullet\bullet\bullet$ - - 84.6 82.6 65.1 49.4
$\bullet\bullet\circ$ $\circ\bullet\bullet$ - 70.9 62.7 27.4 13.7
$\bullet\bullet\circ$ $\bullet\bullet\bullet$ - 85.3 83.4 60.7 45.6
$\circ\bullet\circ$ $\bullet\bullet\circ$ $\circ\bullet\bullet$ 76.9 70.1 31.1 16.0
$\circ\bullet\circ$ $\bullet\bullet\circ$ $\bullet\bullet\bullet$ 83.2 80.3 57.1 41.2
As most recent MLLMs start from pre-trained unimodal backbones, how to align the multimodal content with the language model plays a critical role in the performance of the resulting MLLMs. For example, a common practice adopted by many MLLMs [LiuLWL23a, TongBWWIAYYMWPF24] is the inclusion of a two-stage training schedule: (i) an alignment stage that trains the randomly initialized MLP projector only, followed by (ii) a fine-tuning stage that trains both the projector and language model while keeping the vision encoder frozen. Other studies (e.g., [Karamcheti0BLKS24]) advocate a single-stage training protocol, suggesting that including the projector training stage is unnecessary. Furthermore, when adapting the LLMs to rich visual content such as long videos, some research [apollo] has shown that a three-stage pipeline, which progressively unfreezes different components in different stages, improves the performance.
In this experiment, we test six possible training configurations, ranging from one-stage to three-stage, to evaluate the influence of different training strategies. As shown in Table˜4, we found that a single-stage fine-tuning yields the best results. Adding more stages generally degrades the model performance. We also found that making all parameters trainable in a single stage is important, especially for 3D object detection. This may indicate that, compared to modern image encoders such as SigLIP [ZhaiM0B23] and DINOv2 [DINOv2], current pre-trained point cloud encoders are not as versatile in supporting downstream tasks yet.
4 Experiments
In this section, we compare SpatialLM to the current state-of-the-art on scene layout estimation (Section˜4.1) and 3D object detection (Section˜4.2), respectively. Following our empirical study, we use Qwen2.5-0.5B [qwen2.5] as the base model and Sonata [Sonata] as the point cloud encoder. We set of resolution at the finest level at $2.5\mathrm{cm}$ and employ a single-stage training on our dataset.
Besides testing on established RGBD benchmarks, we also investigate whether our model can handle inputs from other sources, such as point clouds reconstructed from RGB videos (Section˜4.3). Other extensions and more technical details can be found in the appendix.
@@ 4.1 Layout Estimation
Layout estimation focuses on predicting architectural elements, i.e., walls, doors, and windows, within an indoor scene. In this experiment, we include results from two representative baselines, namely RoomFormer [YueKSE23] and SceneScript [SceneScript]. As the state-of-the-art specialist model for layout estimation, RoomFormer employs a highly specialized network architecture, using two-level queries to predict room polygons and their corners, respectively. SceneScript, on the other hand, casts 3D structure estimation as an auto-regressive “next-token-prediction” problem. However, unlike our method, it uses specialized structured language commands to describe the scene and trains a Transformer decoder from scratch.
Following the setting of RoomFormer, we perform evaluation on the Structured3D benchmark [Structured3D], which contains 3,500 residential houses with diverse floorplans. We use the original data split of 3000/250/250 for training/validation/testing, respectively. For RoomFormer, we directly download the model checkpoint from its website [^2] https://github.com/ywyue/RoomFormer for inference. Note that RoomFormer takes in 2D density maps projected from point clouds and outputs 2D layouts, which are converted to 3D by extruding based on the lowest/highest points in the point cloud. For SceneScript, we download the pre-trained model and code [^3] https://github.com/facebookresearch/scenescript, and fine-tune it on Structured3D with the default hyperparameter settings.
@@
RoomFormer [YueKSE23] 42.2 83.4 81.4
SceneScript [SceneScript] 29.9 90.4 89.2
SpatialLM (ft. Structured3D) 603.5 32.6 18.1
SpatialLM (ft. Ours) 603.5 59.9 44.7
SpatialLM (ft. Ours $\rightarrow$ Structured3D) 603.5 94.3 93.5
For our model, we include three variants: (i) trained on Structured3D only, (ii) trained on our dataset only, and (iii) first trained on our dataset, then fine-tuned on Structured3D. As can be seen in Table˜5, directly training LLMs on the smaller Structured3D dataset leads to poor performance. In contrast, by first training on our new dataset and then on Structured3D, our model outperforms both baselines for this task. This demonstrates the benefits of our new dataset.
@@
Ground truth RoomFormer [YueKSE23] SceneScript [SceneScript] SpatialLM
@@
Figure˜4 provides some qualitative comparisons. Since RoomFormer detects room elements independently, we can see that some doors and windows are not attached to any wall. Auto-regressive methods like SceneScript and our method can preserve the relationships between structures by design. But we observe that SceneScript tends to miss some elements in its output.
@@
Next, we compare SpatialLM to existing methods on 3D object detection. We include results from two representative baselines, namely V-DETR [0001GYLLW00G24] and SceneScript. As the state-of-the-art specialist model for the task, V-DETR builds upon the DETR framework [DETR] and makes several task-specific improvements, including a 3D vertex relative position encoding (3DV-RPE), object-based normalization, among others. Following prior practice, we perform evaluation on ScanNet [ScanNet], which contains 1,513 3D indoor scans with annotations of 18 object categories. The training set is composed of 1,201 scenes, while 312 scenes are used for validation/testing. Since auto-regressive models (i.e., SpatialLM and SceneScript) do not produce confidence scores, we report F1 scores instead of mean Average Precision (mAP), as suggested in [SceneScript].
@@
V-DETR [0001GYLLW00G24] 79.4 65.1 56.8
SceneScript [SceneScript] 29.9 49.1 36.8
SpatialLM (ft. ScanNet) 603.5 2.9 0.7
SpatialLM (ft. Ours) 603.5 33.8 22.6
SpatialLM (ft. Ours $\rightarrow$ ScanNet) 603.5 65.6 52.6
For V-DETR, we directly use the model checkpoint downloaded from its website [^4] https://github.com/V-DETR/V-DETR. A confidence threshold of 0.5 is applied for both objectness and semantic predictions, followed by non-maximum suppression (NMS). For SceneScript, we fine-tune the model on ScanNet with the default hyperparameter settings. During fine-tuning, we apply the same data augmentations as used in PTv3 [PTv3] and V-DETR [0001GYLLW00G24].
For our model, we again include three variants: (i) trained on ScanNet only, (ii) trained on our dataset only, and (iii) first trained on our dataset, then fine-tuned on ScanNet. As can be seen in Table˜6, the first variant has very low F1 scores, suggesting that ScanNet itself is too small to train LLMs on. By first training on our new dataset and then on ScanNet, our model outperforms SceneScript and achieves competitive results against the state-of-the-art specialist V-DETR.
Figure˜5 shows some qualitative results. Compared to SceneScript, SpatialLM demonstrates improved detection accuracies across categories. In comparison to V-DETR, we observe that our method has some difficulty in locating objects in “picture”, “sink” and “showercurtain” classes. Notably, “picture” and “sink” are the smallest objects in ScanNet by volume ($0.1\mathrm{m^{3}}$), whereas “showercurtain” category is not included in our dataset and has the fewest occurrences in the ScanNet.
@@
Ground truth V-DETR [0001GYLLW00G24] SceneScript [SceneScript] SpatialLM
@@ 4.3 Zero-shot Detection on Videos
Although significant progress has been made in SFM and SLAM techniques over the past decades, reliably inferring 3D structure from videos remains a highly challenging problem. Recently, DUSt3R [DUSt3R] and its extensions [MASt3R, MASt3R-SfM, MAST3RSLAM, slam3r] provide a novel solution that employs deep networks for simultaneous point matching and camera motion estimation. These techniques are shown to be remarkably robust in dealing with real-world videos or unconstrained image collections.
In this experiment, we collect a set of 107 videos of virtual indoor scene tours and use MASt3R-SLAM [MAST3RSLAM] to reconstruct a point cloud from each video. Then, we test SpatialLM in a zero-shot setting on the reconstructed point clouds. Unlike RGBD scans, these reconstructions have significant noises, occlusions, and geometric artifacts, posing a greater challenge for structured 3D understanding.
Figure˜6 shows some qualitative results. Without fine-tuning, SpatialLM demonstrates strong resilience to imperfect data and preserves consistent layout and object-level predictions. Trained on synthetic datasets with accurate bounding boxes, the model infers full object sizes and orientations from sparse or occluded inputs. In the first example, it predicts beds and nightstands extending to the floor, even when those regions are not visible. Leveraging the auto-regressive nature of LLMs, SpatialLM is also capable of generating plausible room layouts from partial observations. In the second example, the model reasonably reconstructs areas such as the balcony and dining space, filling in missing elements based on contextual understanding.
@@ 5 Conclusion
In this study, we demonstrate the feasibility of training LLMs for structured indoor modeling tasks. We regard it as a meaningful step towards building future foundation models that can not only understand, but also reason about, interact with, and even create structured 3D scenes.
Limitations. SpatialLM has several limitations. First, it falls short of delivering a universal, state-of-the-art model for extracting 3D structure from arbitrary point clouds. There are significant differences between point clouds from different sources, such as monocular videos, RGBD scans, and LiDAR sensors. As shown in Section˜4, although our model exhibits reasonable generalizability across datasets, fine-tuning on a specific dataset is still needed to achieve the best performance. Further scaling up the data and model size could be a promising future direction. Second, in this paper, we focus on training LLM for structured indoor modeling only, without considering how it affects the model’s natural skills in natural language processing and reasoning. To address this, a more thorough benchmarking may be performed. Third, our current approach models indoor layouts using a set of predefined object categories, limiting the LLMs in leveraging their open-ended language capabilities. Future directions include extending our work to support open-vocabulary object detection and 3D visual question answering (VQA), enabling more flexible and generalizable scene understanding.
@@ Appendix A SpatialLM Dataset
Object categories. We have curated a set of 59 commonly occurring object categories, organized by functional and semantic similarity, as shown in Table˜7. Statistics of these categories can be found in Figure˜7.
@@
Figure˜8 plots the object-to-room correlation heatmap, which reveals strong correlations between object distribution and specific room types. Additionally, each room type contains a variety of object categories, highlighting the diversity of our dataset.
@@
Dataset visualizations. We provide visualizations of the ground-truth point cloud and layout annotations in Figure˜9 to showcase the quality of the photo-realistic scans and the realism of the layout. Additionally, we show rendered images of objects in Figure˜10 to illustrate the diversity and quality of the objects in our dataset.
@@ Appendix B Implementation Details of SpatialLM
In Figure˜12, we show the full prompt and an example response of SpatialLM. Note that we shift the point cloud and the corresponding layout and object box coordinates to ensure that they are all non-negative. Next, we quantize the coordinates into 1,280 bins, each with a resolution of $2.5\mathrm{cm}$. LLM predicts integer values, which are mapped back to continuous values using the inverse transformation and normalization, restoring them to the original coordinate system.
Figure 11: Five-level hierarchical structure of the point cloud encoder [Sonata]. $D$ indicates the feature dimension at each level, and grid pool downscales the point cloud resolution by 2 times along each dimension.
Data augmentations. We apply a number of augmentations to the point cloud, which are summarized in Table˜8 and explained below.
We begin with a cuboid crop following V-DETR [0001GYLLW00G24]. This process involves randomly selecting a point as the center of the cuboid and then determining random dimensions for the cuboid relative to the input point cloud size. We maintain a minimum aspect ratio of 0.8 and ensure that at least 50,000 points remain after the cropping process.
@@
cuboid crop min_points=50,000, aspect=0.8, min_crop=0.75, max_crop=1.0, pp=0.1
random jitter $\sigma$=0.025, clip=0.05, ratio=0.8, pp=0.9
random jitter $\sigma$=0.2, clip=0.2, ratio=0.05, pp=0.8
random jitter $\sigma$=0.4, clip=1.0, ratio=0.001, pp=0.75
random jitter $\sigma$=0.5, clip=4.0, ratio=0.0005, pp=0.65
elastic distort params=[[0.2, 0.4], [0.8, 1.6]], pp=[0.8, 0.5]
@@
auto contrast pp=0.2
chromatic translation pp=0.75, ratio=0.1
chromatic jitter std=0.05; pp=0.8
color drop pp=0.1
@@
For experiments on Structured3D [Structured3D] and ScanNet [ScanNet], we use a batch size of 8. Fine-tuning on Structured3D involves 50 epochs, while keeping the other settings the same. For fine-tuning on ScanNet, we introduce an additional pre-training stage to mitigate the difference in label spaces across datasets. Specifically, we map the classes in our dataset to those of ScanNet, and further train the model on our dataset for 3 epochs, followed by fine-tuning on ScanNet for 30 epochs.